OpenStudy (anonymous):

The numerator and the denominator of a fraction are thirteen and fifty, respectively. When a certain number is added to this numerator and subtracted from this denominator, the resulting fraction has a numerator to denominator ratio of two to one. What is the number? 29 37 39

OpenStudy (anonymous):

@surd

OpenStudy (anonymous):

@sourwing

OpenStudy (anonymous):

@surjithayer

OpenStudy (anonymous):

@skullpatrol

OpenStudy (anonymous):

It would be 29. 29+13=42 50-29=21 This would make the ratio two - one.

OpenStudy (anonymous):

@Firejay5

OpenStudy (anonymous):

@eliassaab

OpenStudy (anonymous):

@april115

OpenStudy (anonymous):

Dude stop it. The question was already answered.

OpenStudy (anonymous):

2.(13+x)/(50-x)=2 13+x=2(50-x) 13+x=100-2x x+2x=100-13 3x=87 x=29 he is right

OpenStudy (anonymous):

@surd

OpenStudy (anonymous):

ah whats happening.........thanks

OpenStudy (anonymous):

any more?

OpenStudy (anonymous):

one more

OpenStudy (anonymous):

what?

OpenStudy (anonymous):

Who gave donmath a medal. He has just been tagging people. He never even tried to answer the question.

OpenStudy (anonymous):

hey dobmath

OpenStudy (anonymous):

how slife

OpenStudy (anonymous):

brb

OpenStudy (anonymous):

I was wondering the same thing.. that happens a lot on here though

OpenStudy (anonymous):

@dirtydan667, do you have any more questions that you need help with.

OpenStudy (texaschic101):

is there another question ?

OpenStudy (anonymous):

how slife

OpenStudy (anonymous):

hey dobmath

OpenStudy (texaschic101):

I am wasting my time here....c ya